OpenStudy (anonymous):

PLEASE HELP ME!!! REALLY NEED HELP!!! The equation x/3 - 12 = -7 is solved in several steps below. For each step, choose the reason that best justifies it.

OpenStudy (anonymous):

x/3 - 12 = -7 x/3 =12 -7 x/3 = 5 x = 5*3 x=15